OpenStudy (anonymous):

Find the solution to the equations. 2x - y = -3 x + y = 0 ( ? ),( ? )

OpenStudy (anonymous):

The easiest way to solve a problem like this is to solve one equation for one of the variables, say x. Then whatever you get x equal to, substitute that into the x in the other equation and solve it.

OpenStudy (anonymous):

2x - y = -3 x + y = 0 y=-x 2x-(-x)=-3 3x=-3 x=-1 y=1

OpenStudy (anonymous):

So x+y = 0 gives you x = -y substituting into the first equation gives 2x-y = -3 2(-y) - y = -3 -3y = -3 y = 1 Then just substitute that y value into the second equation and solve for x. x+y = 0 x + 1 = 0 x = -1. So x = -1, y = 1 is your solution.

OpenStudy (anonymous):

thank you!